OpenStudy (anonymous):

Polynomial Operations... SIMPLIFY: (4y^2-5y) + (3y-7y^2) - (2y^2+6y-5)

OpenStudy (anonymous):

A.)9y^2-8y-5 C.)-3y^2+2y+5 B.)-5y^2-8y+5 D.)5y^2-6y-5

jimthompson5910 (jim_thompson5910):

Any ideas?

jimthompson5910 (jim_thompson5910):

Not guesses at the answer, but any ideas on how to tackle this problem?

OpenStudy (anonymous):

\[(4y ^{2}-5)+(3y-7y ^{2})-(2y ^{2}+6y-5)\] \[=4y ^{2}-5y+3y-7y ^{2}-2y ^{2}-6y+5\] \[=4y ^{2}-7y ^{2}-2y ^{2}-5y+3y-6y+5\] \[=-5y ^{2}-8y+5\]